### v2

# VIP: Versatile Image Outpainting Empowered by Multimodal Large Language Model

Abstract.
In this paper, we focus on resolving the problem of image outpainting, which aims to extrapolate the surrounding parts given the center contents of an image. Although recent works have achieved promising performance, the lack of versatility and customization hinders their practical applications in broader scenarios. Therefore, this work presents a novel image outpainting framework that is capable of customizing the results according to the requirement of users. First of all, we take advantage of a Multimodal Large Language Model (MLLM) that automatically extracts and organizes the corresponding textual descriptions of the masked and unmasked part of a given image. Accordingly, the obtained text prompts are introduced to endow our model with the capacity to customize the outpainting results. In addition, a special Cross-Attention module, namely Center-Total-Surrounding (CTS), is elaborately designed to enhance further the the interaction between specific space regions of the image and corresponding parts of the text prompts. Note that unlike most existing methods, our approach is very resource-efficient since it is just slightly fine-tuned on the off-the-shelf stable diffusion (SD) model rather than being trained from scratch. Finally, the experimental results on three commonly used datasets, i.e. Scenery, Building, and WikiArt, demonstrate our model significantly surpasses the SoTA methods. Moreover, versatile outpainting results are listed to show its customized ability. Our source code is available at: https://github.com/ucasyjz/VIP.
Image Outpainting, Large Language Model, Prompt Learning
ccs: Computing methodologies Computer vision problems
Figure 1. The visualization of versatile outpainting results generated by our proposed method. Unlike existing methods (left), our method (right) can generate satisfactory outpainting results in both unconditional and conditional paradigms.
1. Introduction
In recent years, with the tremendous progress in image synthesis field (Yi et al., 2023; Pan et al., 2023; Lu et al., 2023; Qiu et al., 2023), image outpainting has become a striking topic. It aims to generate new content based on the original sub-region of a given image. The extrapolated content of the surrounding regions is expected to be consistent with the original content of the center region. Compared to the image inpainting task (Pathak et al., 2016; Elharrouss et al., 2020; Wang et al., 2023; Yang et al., 2017), outpainting is more challenging due to the lack of adequate contextual information between the extrapolated image region and the original center image region.
In the early stage, a line of methods is devoted to improving image outpainting via structure modification of the base model. For example, SRN (Wang et al., 2019) introduces semantic content re-generation based on the Convolutional Neural Network (CNN) architecture. Then, U-Transformer (Gao et al., 2023) is proposed by replacing the CNN structure with the transformer structure to obtain better results. To address the issue of object outlines in extrapolated regions not aligning with the input sub-images, (Gao et al., 2024) exploit eural Ordinary Differential Equations (Neural ODEs) to build a unique bottleneck architecture that enables seamless extrapolation. With its prevalence in numerous visual generation tasks, there have been recent works attempting to introduce diffusion models (Ho et al., 2020) to resolve the image outpainting problem. PQdiff (Zhang et al., 2024b) learns the correlation between different image regions using the diffusion pipeline.
Although previous works are built based on various architectures, the commonality of them is they all assume the given sub-image should be the sole input for image outpainting as query condition. We consider this assumtion severely limits their application and scalability in real scenarios. Firstly, in comparison to the single-variety output, it is more convenient to generate multiple outpainting results according to the preference of user. Second, most existing approaches are trained from scratch tailored for the outpainting task, which have not fully unleashed the potential capacity of the off-the-shelf large generation models, such as SD (Rombach et al., 2022) and DALLE (Ramesh et al., 2021). By contrast, the paradigm of fine-tuning the large generation models for specific tasks has been proven effective in plentiful related works (Kawar et al., 2023; Karras et al., 2023; Everaert et al., 2023; Zhang et al., 2023a; Moon et al., 2022).
To resolve this problem, we present to introduce the text prompts as semantic guidance of the image outpainting pipeline. Doing so allows us to customize the outpainting results in accordance with our willingness. Moreover, the intervention of text prompts facilitates our exploitation of the well-pretrained text-to-image generation model (Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022).
Besides, in existing works (Zhang et al., 2024b; Yao et al., 2022) for image outpainting, the relation modeling between different image blocks is conducted in random regions. Actually, we argue there is a nonnegligible distinction between the central and surrounding regions of an image. Specifically, in the center regions, the content tends to be noteworthy objects or primary information, while in surrounding regions, the content tends to be background information. Considering image outpainting extends from the central regions to the surrounding regions, it is imperative to capture the semantic association between these two spatial positions explicitly.
This paper addresses the above issues by proposing a novel framework namely Versatile Image Outpainting (VIP). First, to enable this model to be aware of the space information through text prompt as query instruction, we need to construct the pairwise prompt-image data. To achieve it, we leverage the Multimodal Large Language model (MLLM) to produce the textual descriptions of the central and surrounding regions of images, followed by organizing the results into a unified format for output. These operations were performed automatically without the need for manual annotation. Based on the obtained image-prompt data pair, to better leverage the capabilities of the pre-trained large model, we construct our image outpainting model modified from a off-the-shelf stable diffusion (SD) model (Rombach et al., 2022). Furthermore, we elaborately design a Center-Total-Surrounding (CTS) cross-attention module to capture the fine-grained correspondences between the central and surrounding regions of the image and the corresponding parts in the text prompt, which enhances the customization and generalization capabilities of our model. Notably, unlike existing methods, our model is computationally efficient, requiring only fine-tuning on a small amount of data to achieve the current best outpainting effects. As illustrated in Fig. 1, we show that the introduction of the novel text prompt format not only helps us achieve diverse customization results, but also significantly improves the quality of the extrapolated images.
In brief, the contributions of this paper can be summarized as:
  • We build a novel diffusion model based image outpainting framework that is characterized by customizing the results according to the requirements of users. By leveraging the Multimodal Large Language model to annotate images with a novel prompt format automatically, we construct our model via fine-tuning the pre-trained stable diffusion model.
  • A novel Center-Total-Surrounding (CTS) cross-attention module is presented to promote the fine-grained interaction between corresponding parts in images and text prompts.
  • The experimental results not only validate our approach substantially outperforms the existing methods, but also can generate diverse results based on customized instructions.
2. Related Work
Figure 2. The overall framework of our proposed learning method. First, we concatenate the raw image latent, masked image latent, and mask as the input of the unet. The output will be calculated as mean-square error (MSE) loss with the added noise in the raw image latent. Meanwhile, the proposed text prompt and mask will be input into the proposed cross-attention module as conditions. ’R’ and ’C’ mean the resize and concatenation operations, respectively.
2.1. Image Outpainting
The task aims to extrapolate the surrounding contents based on the center contents of the image. The research (Sabini and Rusak, 2018) introduces the deep learning method into the image outpainting task but only performs the horizontal outpainting. IOH (Van Hoorick, 2019) is proposed by utilizing the Generative Adversarial Network (GAN) (Creswell et al., 2018) to implement the image outpainting. SRN (Wang et al., 2019) is developed by directly learning the semantic features for outpainting. Kim et al. (Kim et al., 2021) transform the outpainting task into the inpainting task by swapping the outer region of the image with its inner region. Lu et al. (Lu et al., 2021) propose a bidirectional content transfer module to generate the intermediate region between two different input images. Gao et al. (Gao et al., 2023) further improve the results with a U-transformer structure and extend the horizontal outpainting to all-side outpainting. QueryOTR (Yao et al., 2022) further adopts the ViT module and transforms this task into a patch-wise sequence-to-sequence autoregression problem. Neural ODE (Gao et al., 2024) is designed by implementing continuous extrapolation in latent space to make the generation results smoother. However, these methods are all trained without text prompt conditions, which suffer from the lack of semantic information guidance. Note that Inout (Cheng et al., 2022) also introduces text as conditions to control the outpainting results. Its text prompt annotation is just represented as one label word. This label is annotated by a semantic segmentation model that only performs close-vocabulary classification on a narrow range of classes. By contrast, our employed MLLM could return the open-world results and automatically organize them into templated answers. Moreover, Inout is built based on GAN architecture rather than diffusion models as ours, which limits its capacity.
2.2. Diffusion Model
The diffusion model (Dhariwal and Nichol, 2021; Avrahami et al., 2022; Zhang et al., 2024a; Bar-Tal et al., 2024; Yang et al., 2023a; Zhou and Tulsiani, 2023) is a generative model for producing high-quality, diverse data, such as images, video, and text. These models generate data by gradually introducing noise into the data and then learning how to reverse the process.
DDPM (Ho et al., 2020) lays the foundation of the diffusion era for the generation task. It generates data by simulating the diffusion process of data with high quality and diversity. To speed up the diffusion process by reducing computation costs, LDM (Rombach et al., 2022) transforms the image space into the latent space through a vae encoder and maps it back to the image space using a vae decoder. Meanwhile, it incorporates the text condition into the pipeline for customized generation. Based on LDM, Stable Diffusion is a commonly-used pretrained large-scale model, which has extensively promoted the development of the generation field. ControlNet (Zhang et al., 2023b) is proposed to add conditional control by copying a trainable version based on stable diffusion. DALLE-2 (Ramesh et al., 2022) combines the clip and diffusion model to obtain better text-to-image results. Dreambooth (Ruiz et al., 2023) achieves subject-driven generation by fine-tuning the pretrained stable diffusion model on just a few images of a subject. PQDiff (Zhang et al., 2024b) is proposed by learning the correlation between randomly cropped views for image outpainting based on a diffusion model. However, it still lacks text conditions.
2.3. Multimodal Large Language model
Multimodal Large Language models are designed to process and understand multiple types of data (such as text, images, video, etc.). These models can learn across different data modes and capture associations between different modes to perform well on a range of cross-modal tasks, including Visual Question Answering (VQA), multimodal translation, and more. Here are some important multimodal Multimodal Large Language models. GPT-4 (Yang et al., 2023b) is an open-source MLLM that can understand the image. With the Reinforcement Learning from Human Feedback (RLHF), the feedback from humans enhances the output results. Llama2 (Touvron et al., 2023) is proposed by FaceBook with some improvements in structure and data training. Through three-stage training, Qwen-VL (Bai et al., 2023) further enhances the visual ability of MLLM by introducing the visual encoder. ERNIE-ViL (Yu et al., 2021) adopts scene graphs of visual scenes to build detailed semantic connections across vision and language.
3. Methods
In this paper, we introduce a novel learning framework based on the diffusion model namely versatile image outpainting (VIP), and the whole pipeline is demonstrated in Fig. 2. The details of each component are described in the following.
3.1. Preliminaries
Diffusion Model Diffusion Models are a class of deep learning models that generate high-quality data by gradually increasing and then decreasing the noise on the data. This model performs well in the fields of image generation, video generation, text generation, etc. The working principle of the diffusion model can be divided into two main stages: the forward and the reverse process. The forward process is demonstrated as follows:
(1) $q_{t}(x_{t}|x_{t-1})=N(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$
Then the reverse process to obtain final results is computed by:
(2) $x_{t-1}=\frac{1}{1-\beta_{t}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{a}_{t}}}\epsilon_{\theta}(x_{t},t))$
where $\overline{a}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. Finally, the diffusion model is trained using the following loss ${L}_{DM}$:
(3) $\mathcal{L}_{DM}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$
Multimodal Large Language model Multimodal Large Language models (MLLM) are advanced artificial intelligence technique that combines the ability to process different types of data. These models gain the ability to understand and generate content across multiple modals by learning large multimodal datasets. They can understand the relationship between text and images, generate text prompts from images, etc. The users should give an input prompt to the MLLM for usage. For example, ’description of the input image’, and the MLLM will become an expert in the image caption field and generate the corresponding description. In this paper, we utilize the GPT-4 (Yang et al., 2023b) model due to its excellent performance.
Figure 3. The architecture of Center-Total-Surrounding Cross-attention module. $\cdot$’, ’$\times$’, and ’+’ mean the element-wise multiplication, tensor multiplication and element-wise addition, respectively.
3.2. Data Generation
In this section, the implementation of introducing the text prompt condition into the outpainting task is described. It is considered the traditional prompt format mainly lacks space information, which can not provide accurate condition information for specific image regions. The task needs to restore the surrounding content and keep it consistent with the center content by distinguishing the concrete object contents in different regions. The model should be aware of the precise information about different regions in the generation process. Therefore, incorporating the space information into the text prompt condition is significant. It is well known that the existing Multimodal Large Language models (MLLM) have a strong ability to understand image contents. Therefore, without taking a lot of manual labeling cost, we utilize the Multimodal Large Language model to reconstruct a novel text prompt format, ’Center:xxx; Surrounding:xxx’ (CS), for fine-tuning the model. From our perspective, better results on conditional and unconditional image outpainting are expected. We redefine the conditional and unconditional generation based on whether the keywords after ’Center’ and ’Surrounding’ exist.
Specifically, we define the task prompt as follows, [Please use three English keywords to describe the center and the surrounding of the image, and output in the following format:‘Center:xxx,xxx,xxx; Surrounding:xxx,xxx,xxx’]. The MLLM takes this prompt as input and auto-regressively captions the different regions of the input image. Then, the data pairs of the enhanced text prompt with space information and the corresponding image are obtained. After generation, we filter the data to reduce the noisy samples and set a part of the data with a prompt ’Center:;Surrounding’ without keywords as the unconditional outpainting, and the other parts with keywords as the condition outpainting. Therefore, customized outpainting can be achieved based on the generation of novel text conditions. Then, the corresponding fine-tuning can be performed using the proposed text prompt format.
3.3. Center-Total-Surrounding Cross-Attention
In this section, we will describe the implementation of the Center-Total-Surrounding (CTS) Cross-Attention module using the proposed text prompt format. To achieve the alignment between different image regions and the corresponding text parts, the text prompt part for a specific image area should be input into the model, independently. Therefore, we propose a CTS Cross-Attention to realize the purpose as shown in Fig. 3. Specifically, the details of the implementation are described below.
$F_{I}$ is the latent features of the image, and $F_{T}$ is the latent features of the text prompt. Under normal circumstances, the cross-attention operation between image and text is calculated as follows:
(4) $Q$ $=W_{Q}\cdot F_{I},K=W_{K}\cdot F_{T},V=W_{V}\cdot F_{T}$
$Y$ $=\text{softmax}\left(\frac{{QK^{T}}}{{\sqrt{d_{k}}}}\right)V$
Where $W_{Q}$, $W_{K}$, and $W_{V}$ are linear layers to map the two features.
However, it can be found that the space information is lacking in the original text prompt, and the different text tokens cannot correspond accurately to different image tokens in the masked image without any guidance. Therefore, the CTS cross-attention is proposed to generate the results of the local cross-attention for better conditions. Specifically, we first split the prompt into two parts given a prompt as ’Center:xxx; Surrounding:xxx’ according to the center and surrounding settings. Then, we develop two additional cross-attention modules to control the generation of the center and surrounding regions.
We define the text token embedding of ’Center:xxx’ and ’Surrounding:xxx’ as $F_{T}^{C}$ and $F_{T}^{S}$, and the mapped features are denoted as $K^{C}$, $V^{C}$ and $K^{S}$, $V^{S}$, respectively. Then the specific cross-attention between the image and different text parts is computed by:
(5) $Y^{C}$ $=\text{softmax}\left(\frac{{Q{K^{C}}^{T}}}{{\sqrt{d_{k}}}}\right)V^{C}$
$Y^{S}$ $=\text{softmax}\left(\frac{{Q{K^{S}}^{T}}}{{\sqrt{d_{k}}}}\right)V^{S}$
Table 1. Comparison results with the state-of-the-art methods. The best and second best results are indicated by boldface and underline, respectively. ’+ copy’ means that after the network generates extrapolated images, the model duplicates the input sub-images into the corresponding central regions of the extrapolated images.
Methods Type Scenery Building Facades WikiArt
[…]
SRN (Wang et al., 2019) GAN-based 47.781 2.981 38.644 3.862 76.749 3.629
NSIPO (Yang et al., 2019) GAN-based 25.977 3.059 30.465 4.153 22.242 5.600
IOH (Van Hoorick, 2019) GAN-based 32.107 2.886 49.481 3.924 40.184 4.835
U-Transformer (Gao et al., 2023) GAN-based 20.575 3.249 30.542 4.189 15.904 6.567
Neural ODE (Gao et al., 2024)+copy GAN-based 19.352 3.998 21.009 5.317 14.828 8.957
QueryOTR(Yao et al., 2022)+copy GAN+MAE 20.366 3.955 22.378 4.978 14.955 7.896
PQdiff (Zhang et al., 2024b) Diffusion-based 29.446 3.849 28.855 4.879 10.545 7.374
PQdiff (Zhang et al., 2024b) +copy Diffusion-based 20.100 3.981 19.133 5.350 7.968 8.605
Ours(GLT) Diffusion-based 26.266 4.606 35.054 4.893 10.027 7.096
Ours(GLT)+copy Diffusion-based 23.442 4.479 27.158 5.021 7.326 7.437
Ours(SFT) Diffusion-based 14.883 4.123 27.358 5.146 3.805 7.251
Ours(SFT)+copy Diffusion-based 10.717 4.074 17.775 5.268 3.010 7.337
The initial weights of the two new attention modules are copied from the original attention module to maintain the strong generation ability of the pretrained model. Meanwhile, based on the input mask information, we can separate the latent features of the center and surrounding regions. Then, the final results are obtained by adding the two parts:
(6) $\hat{Y}^{C}$ $=Y^{C}*(1-mask)$
$\hat{Y}^{S}$ $=Y^{S}*mask$
$\hat{Y}$ $=\hat{Y}^{C}+\hat{Y}^{S}$
Then we retain the latent features $Y$ obtained by performing cross-attention with the total text prompt to maintain the global consistency. Here, we introduce a trainable parameter $a$ to fuse the two latent features adaptively:
(7) $Y$ $=(1-a)*Y+a*\hat{Y}$
Therefore, local conditions for specific regions and global consistency across the whole image can be accomplished. Customized outpainting results are obtained with our proposed method.
4. Experiment
4.1. Dataset
Laion2B-en (Schuhmann et al., 2021, 2022) The dataset contains about 2B image-text data pairs. We apply the data generation strategy to generate 224 pieces of data with the proposed text prompt format.
Scenery (Yang et al., 2019) The Scenery dataset consists of many images of natural landscapes and has about 5,000 images for training and 1,000 images for testing.
Building Facades (Gao et al., 2023) The images in this dataset are focused on urban buildings. There are about 16,000 images in the training set and 1,500 images in the testing set.
WikiArt (Tan et al., 2016) WikiArt is a large dataset of paintings by 195 artists, which contains about 45,500 training images and 19,500 testing images. It has different degrees of realism and stylization.
Figure 4. The visualization comparison of the outpainting results generated by different methods. Our proposed approach surpasses other well-behaved methods in the concerned box region.
[…]
A woman and a girl in white dresses sit in an open car.
4.2. Implementation Details
In this section, the implementation details are described briefly. Based on the pretrained stable-diffusion-inpainting model, we perform two training processes. The one, denoted specific training (SFT), is fine-tuning the model on Scenery, Building, and Wikiart dataset for comparison of the SoTA results with the existing methods in the proposed unconditional type, ’Center:; Surrounding:’, respectively. As existing methods do, the input size is set to $192\times 192$. Unlike the traditional hundreds of epochs for training, the fine-tuning iteration is set to only 30000 with a mini-batch 4. The original images from the datasets are first resized to the set size as ground-truth images. The contents of the region with a size $128\times 128$ in the center are preserved, and the surrounding region is masked off. To show the generalization ability of the proposed method and achieve customized results, the other, namely generalized training (GLT), is fine-tuning the model on the data generated by MLLM from Laion2B-en dataset. With an input size of $256\times 256$, the fine-tuning iteration is set to 10000 with a mini-batch 4. A common optimizer, Adam (Kingma and Ba, 2014), is selected with a learning rate of 5e-6. In the inference stage, the center size and the extrapolation ratio are set to the same as the corresponding training process.
For evaluation, the commonly used Frechet Inception distance (FID) (Heusel et al., 2017) and Inception Score (IS) (Salimans et al., 2016) are chosen as the indicators to show the quantitative performance.
4.3. Compared with SoTA methods
In this section, we compare the outpainting results obtained by the proposed method with those obtained by existing SoTA methods, and the superiority of the proposed method is proven. For quantitative comparison, we select the methods, SRN (Wang et al., 2019), NSIPO (Yang et al., 2019), IOH (Van Hoorick, 2019) , U-Transformer (Gao et al., 2023), Neural ODE (Gao et al., 2024), QueryOTR(Yao et al., 2022) and PQDiff (Zhang et al., 2024b). The statistical data of FID and IS indicators are recorded in Table 1.
As shown in the third and fourth lines from the bottom, performing generalized training (GLT) can achieve a comparable result with the well-behaved methods, indicating that our proposed method has a strong generalization ability and can extend to the outpainting tasks on various image types with more realistic and consistent results. In addition, after performing specific training (SFT) on the three datasets, the results far exceed those of existing methods, as shown in the last two lines. The FID indicator achieves the best level at 10.717, 17.775, and 3.010 on the Scenery, Building, and Wikiart datasets, respectively. Meanwhile, the IS indicator also achieves excellent and advanced results. Especially it can be seen that we do not need to perform fine-tuning on the wikiart dataset and can achieve the SoTA FID results just based on the fine-tuning model on the generated data from the Laion-2B dataset.
Table 2. Results of using each component in the proposed method
        Method         FID $\downarrow$         IS $\uparrow$
        Baseline         178.58         4.819
        CS         16.827         4.031
        CS + CTS         14.883         4.123
In addition, the leading results are achieved by just fine-tuning only one-digit epochs instead of fine-tuning the model for a hundred epochs as existing models, which greatly saves training costs based on the strong generation ability of the pretrained model. Meanwhile, it can be found that whether or not replacing the center content of the generated results with the original center content, our results both exceed the existing best FID result. This means our method can sense the spatial position relation and generate the content of the center and surrounding regions precisely. Moreover, the FID boost effect is not as obvious as the previous method after replacing the center content, which also proves our method can generate whole images with consistent and continuous content.
We also compare the visualization results with other methods. For a fair comparison, we select the same image cases as research (Gao et al., 2024) as shown in Fig 11. Inside the selected box regions, it can be found that the texture details are more clearly visible in the surrounding regions. For example, in the first line, the edge of the dune is smoother and clearer than that of other methods. The outpainting performance has been greatly improved in terms of overall consistency and continuity.
4.4. Ablation Study
In this section, we analyze the effectiveness of each component in the proposed learning method through some ablation experiments on the test set of the Scenery dataset (Yang et al., 2019). We set the outpainting ratio to 2.25x, like most works. The size of the input center region is 128x128, and the size of the output extrapolation image is 192x192. Both qualitative and quantitative results are provided below.
Figure 5. The visualization comparison of the ablation results of each component in the proposed method.
[…]
A woman and a girl in white dresses sit in an open car.
Table 3. Results of using different unconditional paradigms
       Method        FID $\downarrow$        IS $\uparrow$
       prompt=”        15.421        4.131
       Ours        14.883        4.123
Components of The Proposed Framework. To show the contribution of each component, the quantitative results are demonstrated in Table 4. It is evident that the contribution of each component to the final success of our proposed method is important. The first line shows results obtained by directly using the pretrained stable-diffusion-inpainting model. The unsatisfactory results demonstrate that the outpainting task is fundamentally different from the inpainting task, and the model optimized for the inpainting task can not be directly used in the outpainting task to obtain a satisfactory result. After fine-tuning the model on the proposed image-text data pair with space information, ’Center:; Surrounding’ (CS), the FID indicator has been greatly improved, as shown in the second line. This means the distribution of extrapolated images is more similar to that of the ground truth images. The IS indicator can also maintain a comparable level. Meanwhile, the effectiveness of CTS cross-attention is also verified by further using it, and the results are shown in the third line. The FID and IS indicators are both improved compared to those obtained by only using the proposed text prompt format.
Figure 6. The visualization comparison of using different unconditional paradigms.
We further analyze the effectiveness of each component through the qualitative results as shown in Fig. 5. As we can see, the baseline can not behave well in the outpainting task without fine-tuning. With the proposed prompt format, the outpainting results are obviously improved. Moreover, adding the CTS cross-attention further improves the texture details and reduces abnormal generations.
vs. Traditional Unconditional Prompt Format. In this part, we compare the fine-tuning results obtained by the traditional unconditional prompt format, prompt = ’, and the proposed unconditional prompt format, ’Center:; Surrounding:’. The results are recorded in Table 3. It can be found that our proposed format can achieve better results in the FID indicator with a negligible decrease in the IS indicator. The visualization of the results is also demonstrated in Fig. 6. It can be found the our method is better than the traditional unconditional format regarding reality and consistency. The performance of the unconditional outpainting results has been improved after utilizing the proposed method.
Figure 7. The visualization comparison of the portrait results.
Figure 8. The visualization comparison of the customized outpainting results with different text prompts in diverse scenes.
4.5. Portrait Results
In this section, we demonstrate the outpainting results for a more difficult scene, portrait generation. The visualization results on the WikiArt (Tan et al., 2016) dataset are provided in Fig. 7. It can be found that the extrapolated parts are seamless with the original portrait, and the body parts are continuous.
4.6. Customized Results
Thanks to the pretrained stable-diffusion-inpainting model, customized outpainting can be more easily achieved based on the strong generation ability with text prompt conditions after fine-tuning. In this section, the fine-tuned model by performing the generalized training (GLT) is utilized for the customized outpainting. Then we select four samples randomly to provide the corresponding visualization results. Based on the proposed method, we can achieve customized outpainting results according to the proposed text prompt condition, as shown in Fig 8 (a).
For regular outpainting, it can be found that the generation results change correspondingly after adding different surrounding keywords. For example, if we set the keyword of the surrounding region to ’shirt’, ’cap’, ’sign’, and ’box’, the generated results correspondingly change to an accurate object. Therefore, precise conditional customized outpainting can be achieved through our proposed method. In addition, we can also utilize the MLLM to imagine the surrounding content and output more abundant keywords instead of an artificial mode. Then, the various outpainting results can be obtained without a tedious human design process.
In addition, as shown in Fig. 8 (b), we also perform the customized outpainting with an irregularly shaped mask to show the generalization ability of our approach. The noteworthy object in the center region often has an irregular shape. Therefore, we utilize the Grounded SAM (Ren et al., 2024) to extract the main objects in the center region and mask the surrounding region. Then we obtain the masked image and the corresponding mask to perform the irregular outpainting task.
For example, it can be found that in the first case, according to elements in the provided prompt, we can continuously add the surrounding content of the airplane with ’desert’, ’runway’, ’mountain’, and ’sky’, and the corresponding content will appear in the generated result. Moreover, as we continue to add keywords, the generated result maintains good consistency and continuity on the basis of the original result. Therefore, our model can not extrapolate the region of common square shape like the existing methods but also extrapolate the region of irregular shape. Our model can adapt to various outpainting scenarios and obtain satisfactory results.
5. Conclusion
In this paper, we propose a novel framework namely VIP to complete the image outpainting task with Multimodal Large Language model. The method can introduce text prompts with the space information of center and surrounding regions to perform precise outpainting. With the help of the Multimodal Large Language model, we design a new text prompt format called ’Center:;Surrounding:’ (CS) to assign the different object components to specific regions. Then, the proposed novel Center-Total-Surrounding (CTS) cross-attention can further enhance the correlation between each image region and the corresponding text part through an adaptive fusion of multiple cross-attention operations. Therefore, the model can be aware of the exclusive characters of different regions and generate consistent results across the whole image. We carry out extensive experiments on widely-used datasets. Leading quantitative and qualitative results are obtained in both unconditional and conditional outpainting tasks. The effectiveness of our method is verified. In future work, we intend to continue to explore a more effective text prompt format for customized outpainting results.
Appendix A Overview of Appendix
In this supplementation material, we provide additional details that were omitted in the main manuscript owing to the limited space. As for implementation details, we first give more description of the experiment environment and settings and the usage of the Center-Total-Surrounding (CTS) cross-attention. Then, more experiments are conducted to show the effectiveness of the proposed approach, containing a comparison with an open-source outpainting method and an analysis on more different hyperparameter settings. Finally, the visualization comparison with other SoTA methods on more examples is demonstrated. Moreover, various outpainting results obtained by our proposed method are provided below.
Appendix B More Implementation Details
B.1. Training
In this paper, we utilize the pretrained stable diffusion inpainting with an input channel of the unet as 9, which has been fine-tuned explicitly for inpainting. In the fine-tuning process, the whole parameters of the unet are optimized, and other parts like the text-encoder and the vae are frozen. Our sampler utilizes the DDIM (Ho et al., 2020) sampler with 50 sampling timesteps. In the training stage, there is only a traditional loss function of calculating the mean-square error (MSE) between the predicted noise and added noise. The experiments are carried out on the NVIDIA V100 with Pytorch 1.13.
B.2. Center-Total-Surrounding Cross-Attention
For Center-Total-Surrounding (CTS) Cross-Attention, we introduce the module in the downsampling, middle, and upsampling processes, where the original cross-attention module exists. After calculating the resnet layer, the three cross-attention modules are calculated and the corresponding results are adaptively fused for the subsequent forward process.
B.3. The examples of The Generated Training Data
In this section, we will demonstrate some examples of the training data generated by MLLM. The selected image-text pair samples are visualized in Fig. 9. In the future, we will release the whole generated data.
Figure 9. Qualitative samples generated by MLLM.
[…]
A woman and a girl in white dresses sit in an open car.
It can be seen that the content keywords of the center and surrounding regions are effectively split in the four generated samples. With our proposed approach, we can make full use of data in this format for better outpainting performance.
Appendix C More Experiments
Figure 10. Customized results compared with the open-source outpainting method, SD-infinity Outpainting.
[…]
A woman and a girl in white dresses sit in an open car.
C.1. Comparison of Customized Results
In the main manuscript, we have made a comparison with the leading methods reported in the published paper. To further show the advantage of our work, we further compare it with a unpublished but open-sourced method  [^1] https://github.com/lkwq007/stablediffusion-infinity, namely SD-infinity Outpainting here, which is also based on the pretrained stable diffusion of the traditional prompt. With the same setting of the sampling parameters, the visualization results are demonstrated in Fig 10.
It can be found that after using the proposed approach, the best outpainting results are obtained, and the proposed text prompt paradigm with just some keywords is relatively simple compared to the traditional prompt paradigm with a complete sentence. In the first line, the dividing line between sand and road is clearer and more rational. In the second line, the cap generated by our approach has a more normal shape and a more continuous color. In the third line, the waist area is more realistic, and there are fewer tattoos on the arms In the last line, the generated lane obtained by the open-source outpainting repository is more disorganized.
Table 4. Results of using different parameter formats
        Method         FID $\downarrow$         IS $\uparrow$
        baseline         15.421         4.131
        RIF         15.170         4.216
        CF ($a=0.5$)         15,270         4.224
        LF         14.883         4.123
C.2. Analysis of Different Hyperparameter Setting
In this section, we analyze the influence of setting the hyperparameter $a$ used in the CTS cross-attention module. Specifically, we propose three variations for comparison with our used strategy: the random initialization format (RIF), the constant format (CF), and the used learnable format (LF). The corresponding results on scenery are recorded in Table 4.
It can be found that no matter which format we use, the FID results obtained are always better than those obtained by the baseline, and the IS indicator maintains an excellent level. This indicates that our method is insensitive to the parameter format. The best FID results are achieved by utilizing the learnable format compared to the other two formats. It is understandable that directly setting a constant or random initialization will not adapt to the fine-tuning process and obtain better results. Moreover, using the random initialization format is better than the constant format, which means that using the constant format needs to perform extensive experiments to explore the best value, which is very consuming. Therefore, the learnable format is a reasonable and effective strategy.
Appendix D More Results
D.1. Traditional Unconditional Results
Figure 11. The demonstration of more samples compared to the other methods on three datasets.
[…]
A woman and a girl in white dresses sit in an open car.
In this section, we provide the visualization of other samples used in research (Gao et al., 2024). The unconditional outpainting results on the Scenery, Building, and Wikiart datasets are demonstrated in Fig. 11. As we can see, the texture details are more pronounced around the edges, and the contents are more continuous and not as abnormal as in other methods. For example, the extrapolated ship in the first line has a more normal shape, and the building in the fourth line has more structural information.
Appendix E More Results
E.1. Traditional Unconditional Results
Figure 12. The demonstration of more outpainting results on the three datasets.
[…]
A woman and a girl in white dresses sit in an open car.
In addition, we provide more outpainting results generated by our approach on the three datasets in Fig. 12. Left, middle, and right are the Scenery, Building Facades, and WikiArt dataset, respectively. For input, only the part in the black box area of the GT image is reserved, and the other parts are extrapolated. It can be found that the extrapolated part is consistent with the input part, and the outpainting performance is satisfactory on various samples.
E.2. Customized Conditional Results
In this part, more conditional customized results obtained on samples from the Laion-2B dataset are also shown in Fig. 13. It can be seen that with various prompts, we can obtain visually plausible customized outpainting results.
Figure 13. The demonstration of more customized outpainting results.
[…]
A woman and a girl in white dresses sit in an open car.